sic Equilibrium Calculations 25 of 33 below the reaction. Rank the extent of each rea refers to how much reactant is converted to product at equilibrium. equilibrium constant for the following two reactions. The equilibrium concentration of each reactant and product is given in the box action at equilibrium by calculating the equilibrium constant. Note that the extent of each reaction Reaction 2 2N2(g) +O2(g) Reaction 1 CO(g) + O2(g) 2CO2(g) 2NO2(g) [CO] = 1.1 M [02] = 1.1 M [CO2-8.8 M [N2-1.8 M 0] 2.0 M [NO2] = 0.02 M Compare Ke values to indicate which reaction proceeds to a greater extent at equilibrium. View Available Hint(s) O Reaction 2, with Ke-1.6 x îof > Reaction 1, with K, 1.6 x 10-2 O Reaction 1, with Kc 727 > Reaction 2, with K, 5.6 × 10.3 O Reaction 2, with K, = 64.0 > Reaction 1, with K, 6.2 x 10-5 O Reaction 1, with Ke-64.0 > Reaction 2, with K, 6.2 ×。5 9
Answer #1

CO(g) + O2(g) -------------> 2CO2(g)

Kc    = [CO2]^2/[CO][O2]

        = (8.8)^2/(1.1)*(1.1)   = 64

2N2(g) + O2(g) -----------> 2NO2(g)

Kc   = [NO2]^2/[N2][O2]^3

         = (0.02)^2/(1.8)^2 *2 = 6.2*10^-5

Reaction1 with Kc = 64.0> Reaction 2 ,with Kc = 6.2*10^-5 >>>>>answer

Plz any doubt comment on the BOX